# LaST0: Latent Spatio-Temporal Chain-of-Thought for Robotic Vision–Language–Action Model

### v1

@@ III-C Latent Spatio-Temporal Chain-of-Thought
By fusing visual information, 3D geometric structure, and robot proprioception, the model develops a holistic understanding of both the physical world and the robot itself. The interleaved multimodal structure further encourages the model to learn the coupled dynamics across different modalities over time. Notably, the temporal granularity can be flexibly adjusted: we either adopt keyframe extraction as in [67] or use densely sampled frames, depending on task requirements. Moreover, by compressing high-dimensional sensory inputs into a latent sequence of length $3\times H$, we avoid the prohibitive cost of decoding pixel-level images or long textual sequences. As discussed in Section III-D, we introduce an asynchronous frequency design for the two experts, which further accelerates action generation.
Sequence Structure. To better organize LaST CoT reasoning and action generation, we introduce three special tokens: <latent_start>, <latent_end>, and a placeholder token <latent_pad>. The reasoning segment is structurally defined as a sequence bounded by the start and end tokens, with the intermediate positions reserved for the latent embeddings. During Training, we replace the intermediate <latent_pad> tokens with the ground-truth latent sequence $\mathcal{Z}_{\text{GT}}$. This allows the model to learn the transition dynamics via standard teacher forcing. While during Inference, the model is initialized with <latent_start> followed by a sequence of <latent_pad> tokens. The slow reason expert then autoregressively generates continuous embeddings, sequentially filling the positions of the <latent_pad> placeholders until the pre-defined horizon is filled. The horizon length can be adaptively adjusted, and its effect is further analyzed through ablation studies.
Latent Supervision Strategy. We train the slow reasoning expert to predict latent CoT embeddings in an autoregressive manner, using continuous latent regression rather than discrete token likelihoods [84, 76]. Specifically, the slow expert is trained to predict a sequence of latent reasoning states $\hat{\mathcal{Z}}$ in a next-step prediction manner, conditioned on preceding observations and context. Unlike conventional CoT supervision based on discrete token prediction, our latent targets consist of continuous, high-dimensional embeddings that encode future physical world states. To align the predicted latent semantics with the ground-truth representations, we employ cosine similarity as the supervision objective. The loss is defined as:
@@
By maximizing directional alignment in the latent space, this objective encourages the model to anticipate future physical dynamics in a structured and compact manner. The slow expert learns a physically grounded latent reasoning trajectory that serves as a stable scaffold for the action generation.
III-D Dual-System Coordination
@@
Figure 3: The slow reasoning expert performs low-frequency latent CoT reasoning to capture long-horizon spatio-temporal dependencies, while the fast acting expert generates actions conditioned on high-frequency observations and periodically updated latent knowledge. Benefiting from training with a mixture of fast–slow operating ratios, the model can flexibly adjust the reasoning–action frequency ratio at test time.
@@
Asynchronous Frequency Coordination. To coordinate LaST CoT reasoning with high-frequency robotic control, we introduce an asynchronous frequency mechanism between the slow reasoning expert and the fast acting expert. As shown in Fig. 3, we decouple their operating frequencies by a fixed update ratio $\kappa$ (e.g., $\kappa\in{2,4,8}$). The slow expert is activated only at sparse keyframes ($t\bmod\kappa=0$), where it performs autoregressive latent CoT reasoning. In contrast, the fast expert runs at the native control frequency and remains active at every timestep. Between consecutive keyframes ($t\bmod\kappa\neq 0$), the slow reasoning expert is held dormant, while the fast expert generates actions by conditioning on the most recent latent reasoning output. This strategy ensures that action generation is continuously guided by stable, temporally consistent latent conditions, while minimizing redundant computations for LaST CoT reasoning.
For the inputs to the two experts, the slow reasoning expert receives the natural language instruction $l$ and the low-frequency observation $I_{\text{slow}}$, constructing the Latent CoT that encapsulates future physical dynamics. Conversely, the fast acting expert is optimized for rapid closed-loop feedback and receives only the high-frequency observation $I_{\text{fast}}$. Crucially, since our MoT architecture maintains a unified token sequence, the fast expert can efficiently attend to both the linguistic goal and the Latent CoT tokens via the shared attention mechanism.
Inference via KV Cache. During inference, the Key-Value states computed by the slow expert (encapsulating the Latent CoT) are cached in memory. During intermediate fast-control steps, the acting expert only encodes the current observation and attends to the frozen latent CoT cache, effectively retrieving the CoT tokens with $O(1)$, without re-invoking the slow reasoning process. This approach eliminates repetitive decoding, allowing our model to achieve an inference speed of 15.4 Hz on a single RTX 4090 GPU when operating with a 1:4 fast–slow frequency ratio.
III-E Training Recipe
Large-Scale Robotic Pretraining. Both the latent reasoning expert and the action expert are initialized from the Janus-Pro 1B foundation model. As shown in Fig. 2 c), Stage 1 performs large-scale pretraining on a diverse corpus of over 400K trajectories aggregated from Open-X-Embodiment [62], DROID [41], ROBOMIND [82, 33], and other robotic datasets. Notably, beyond following prior VLA works in applying data quality filtering [42, 52], we additionally ensure that all robot state annotations are accurate and physically consistent. Detailed statistics and descriptions of the pretraining data are provided in Appendix -A. This stage optimizes the model on broad robotic datasets to establish a shared representation space, enabling seamless interaction between reasoning and execution within the unified VLA framework.
@@ III-E Training Recipe
Supervised Fine-Tuning (SFT). As shown in Fig. 2, we employ a decoupled 2-stage SFT in downstream adaptation, aiming to prevent the catastrophic forgetting of reasoning capabilities when training action expert. In stage 2, we first optimize the slow reasoning expert while freezing the action expert. By exclusively optimizing the Latent CoT regression loss $\mathcal{L}_{\text{latent}}$, the slow expert aligns its latent representations with domain-specific physical dynamics. Subsequently, in stage 3, we freeze the converged slow expert and load the robotic data pretrained weights for the fast expert. The fast acting expert is subsequently fine-tuned using the standard Flow Matching loss $\mathcal{L}_{\text{flow}}$. While the loss formulation and noise injection scheme follow $\pi_{0}$ [5], our approach differs in that the entire LLM is directly used as the action denoising model, without introducing any additional action heads. Meanwhile, the action expert is trained with randomly mixed fast–slow operating ratios (e.g., 1:1, 1:2, 1:4), which exposes it to latent conditions updated at varying delays. As a result, LaST0 can flexibly accommodate different reasoning-update frequencies at deployment and adaptively choose the fast–slow inference rate. In ablation studies, we find that training with mixed fast–slow operating ratios does not degrade performance; instead, it improves the model’s robustness during inference.
IV Experiment
In Section IV-A, we evaluate the manipulation performance and inference efficiency of LaST0 against prior methods in simulated environments. The contribution of each component is analyzed through ablation studies in Section IV-B. Section IV-C presents both quantitative and qualitative results of LaST0 on real-world manipulation tasks, including single-arm and dual-arm control under diverse robot configurations.
@@ IV Experiment
TABLE I: Comparison of LaST0 and baselines on RLBench. All methods are trained in the multi-task setting [67], and we report mean success rates (S.R.). Inference speed is evaluated on an NVIDIA 4090 GPU.
@@
box laptop lid seat down to dustpan fridge on base out off hanger rack plants & Var $\downarrow$ speed $\downarrow$
OpenVLA [42] 0.60 0.35 0.75 0.55 0.85 0.20 0.30 0.15 0.20 0.05 0.40 $\pm$0.02 6.3 Hz
SpatialVLA [64] 0.80 0.70 0.85 0.20 0.80 0.15 0.25 0.40 0.15 0.30 0.46 $\pm$0.03 7.9 Hz
CogACT [47] 0.90 0.80 0.95 0.50 0.85 0.50 0.55 0.45 0.30 0.25 0.61 $\pm 0.04$ 9.8 Hz
CoT-VLA [92] 0.95 0.75 1.00 0.80 0.65 0.50 0.40 0.50 0.55 0.50 0.66 $\pm 0.03$ 1.1 Hz
$\pi_{0.5}$ [35] 0.90 0.95 0.85 0.75 1.00 0.05 0.10 0.80 0.75 0.35 0.65 $\pm$0.04 13.8 Hz
HybridVLA [50] 0.85 0.95 1.00 0.90 1.00 0.50 0.50 0.70 0.50 0.50 0.74 $\pm$0.04 6.1 Hz
LaST0 0.95 0.95 1.00 0.80 0.85 0.75 0.75 0.70 0.85 0.65 0.82 $\pm$0.03 15.4 Hz
Figure 4: Ablation study on key design choices of LaST0. We analyze (a) the importance of different latent modalities, (b) the number of tokens allocated per latent modality, (c) the temporal coverage in latent reasoning, and (d) the collaboration frequency between reasoning and action experts. Results are reported as average success rates across 10 RLBench tasks, demonstrating the contribution of each component to overall performance.
IV-A Simulation Experiment
Data collection. We evaluate LaST0 on a diverse set of 10 manipulation tasks from the RLBench [36] benchmark, conducted in the CoppeliaSim simulation environment. The evaluated tasks include 1) Close box, 2) Close Laptop, 3) Toilet seat down, 4) Sweep to dustpan, 5) Close fridge, 6) Phone on base, 7) Take umbrella out, 8) Frame off hanger, 9) Wine at rack, and 10) Water plants. All tasks are executed using a Franka Panda robotic arm with a single front-view observation. Demonstration data are collected by following pre-defined waypoints with motion plans generated via the Open Motion Planning Library [68]. Following the frame-sampling protocol adopted in prior work [67, 27, 38], we construct a training dataset comprising 100 demonstration trajectories per task.
Training and evaluation protocol. We benchmark LaST0 against six representative state-of-the-art VLA models: OpenVLA [42], $\pi_{0.5}$ [35], CogACT [47], SpatialVLA [64], CoT-VLA [92] and HybridVLA [50]. For all baselines, we initialize from the officially released pretrained checkpoints and follow the full fine-tuning configurations recommended by the paper. For CoT-VLA, we reimplement the method on top of Janus-Pro [12], the same foundation model used in our approach, and reproduce its explicit CoT reasoning to ensure a fair comparison. For LaST0, each observation consists of a single RGB image resized to $384\times 384$, a point cloud uniformly subsampled to 1024 points, a language instruction obtained directly from the simulator, and the robot proprioceptive state synchronized with the predicted action. Note that point cloud data are used exclusively during training to construct latent ground truth and are not needed during inference. The model is trained for 40 epochs in Stage 2 and 300 epochs in Stage 3, using the AdamW optimizer [53] across 8 NVIDIA A800 GPUs. Evaluation follows the protocol of prior work [47, 27]. For each task, we perform 20 rollout trials using the final checkpoint, repeat the evaluation three times with different random seeds, and report the mean success rate together with its variance.
Quantitative analysis. As summarized in Table I, LaST0 achieves a mean success rate of 82% across 10 RLBench manipulation tasks, substantially outperforming all prior baselines. In particular, LaST0 surpasses the strongest existing methods, HybridVLA-7B (74%), $\pi_{0.5}$-3B (65%), and CogACT-7B (61%), by margins of 8%, 17%, and 21%, respectively. Beyond the overall average, LaST0 attains the highest success rate on 7 out of 10 tasks, indicating consistent performance gains across diverse manipulation skills, while also showing relatively low variance ($\pm$0.03), reflecting stable execution across tasks. These improvements primarily stem from the latent reasoning representations produced by the reasoning expert, which condition the action expert with compact latent states encoding future visual dynamics, 3D spatial structure, and robot proprioceptive information, enabling more stable and temporally coherent action generation. In terms of execution efficiency, LaST0 operates at an inference speed of 15.4 Hz with an action chunk size of one, which is significantly faster than explicit CoT methods (CoT-VLA: 1.1 Hz) and remains competitive with $\pi_{0.5}$ (13.8 Hz). This demonstrates that LaST0 achieves superior task performance without sacrificing real-time control efficiency. Moreover, as shown in Fig. 5, we compare the attention heatmaps of LaST0 against variants without CoT and with explicit CoT (CoT-VLA). While the no-CoT variant and the explicit CoT method fail to aggregate features on the manipulated objects and ignore the relationship between the robot and the manipulated objects, LaST0 exhibits a highly concentrated attention pattern, highlighting its superior understanding of spatio-temporal information in downstream tasks.
TABLE II: Comparison of LaST0 and baselines in real-world manipulation tasks. We report success rates (S.R.) for standard single-arm and dual-arm tasks (left), and a long-horizon task evaluated over three consecutive steps (right).
Models Wipe Press Place dish Place egg Scoop popcorn Open pot Mean Place egg on bread (Long-horizon)
blackboard stamp on rack on bread into bowl pick corn S.R. $\uparrow$ Step 1  $\rightarrow$ Step 2    $\rightarrow$ Step 3
SpatialVLA [64] 0.60 0.67 0.30 0.20 0.27 0.40 0.39 0.20 0.07 0.00
$\pi_{0.5}$ [35] 0.60 0.73 0.60 0.47 0.53 0.60 0.59 0.47 0.20 0.07
CoT-VLA [92] 0.53 0.60 0.66 0.53 0.33 0.53 0.53 0.33 0.13 0.07
LaST0 0.73 0.93 0.80 0.66 0.66 0.53 0.72 0.66 0.47 0.33
IV-B Ablation Study
To validate the key design choices of LaST0, we conduct comprehensive ablation experiments on 10 RLBench tasks, using the same training and evaluation settings as in the simulation experiments. The results are summarized as follows.
Importance of latent CoT modalities. As shown in Fig. 4 a), we analyze the contribution of each latent modality by selectively ablating individual latent components while keeping the overall architecture unchanged. When using only the image, point cloud, or robot state latent, the model achieves success rates of 74%, 76%, and 75%, respectively, indicating that each modality-specific latent provides a strong basis for action generation. The combination of multiple modality latents continues to provide additional performance improvements, even when the manipulation accuracy is already high. These results validate the importance of modeling comprehensive physical dynamics in the latent space, and further demonstrate that enabling the model to autonomously reason about the relationship between the robot and its interactive environment is effective for robotic manipulation.
@@ IV-B Ablation Study
Number of tokens per latent modality. As shown in Fig. 4 b), we study how the token budget allocated to each latent modality affects performance by varying the number of latent tokens while keeping other components fixed. When no latent token is used, performance drops substantially to 68%, indicating that the absence of latent representations severely limits the model’s reasoning capacity. Introducing a single token for each modality leads to a sharp improvement (up to 82%), demonstrating that even a minimal latent representation is sufficient to establish an effective latent decision state.
Further increasing the number of tokens yields no significant accuracy improvement, suggesting that high-level latent tokens can compactly encode physical information sufficient for effective reasoning.
@@
Temporal coverage in latent reasoning. As shown in Fig. 4 c), we investigate the effect of the temporal horizon used in latent reasoning by varying the number of future time steps encoded into the latent state. Performance improves consistently as the latent temporal coverage increases (from 68% to 82% when extending from 0 to 4 steps), indicating that incorporating longer temporal dependencies enables more informed latent decision states. At the same time, expanding the temporal horizon introduces higher inference latency, revealing a trade-off between decision quality and computational efficiency in reasoning expert. Due to our fast–slow system design, extending the temporal horizon of the latent space does not significantly affect action generation speed.
@@
Collaboration frequency between reasoning and action experts. As shown in Fig. 4 d), varying the collaboration frequency has a clear impact on task success. Fixed ratios such as 1:1, 1:2, and 1:4 achieve comparable performance (75-79%), while overly infrequent collaboration (1:8) leads to a relative drop to 74%. Finally, the mixed strategy trained by combining data from all four collaboration ratios achieves the best performance (82%), using a 1:4 ratio during testing. These results indicate that our frequency joint training scheme enables more robust coordination between the reasoning and action experts across tasks.
Figure 6: Real-world Experiment Setup. We show the details about robot-arm setup and assets of manipulation tasks.
Figure 7: Visualization of task execution processes by real-world single-arm and dual-arm robots (from left to right).
IV-C Real-World Experiment

### v3

@@ 3.3 Latent Spatio-Temporal Chain-of-Thought
The interleaved multimodal structure further encourages the model to learn the coupled dynamics across different modalities over time. Notably, the temporal granularity can be flexibly adjusted: we either adopt keyframe extraction as in (Shridhar et al., 2022) or use densely sampled frames, depending on task requirements. Moreover, by compressing high-dimensional sensory inputs into a latent sequence of length $3\times H$, we avoid the prohibitive cost of decoding pixel-level images or long textual sequences. In Section 3.4, we introduce an asynchronous frequency design for the two experts, which further accelerates action generation.
Sequence Structure. To better organize LaST CoT reasoning and action generation, we introduce three special tokens: <latent_start>, <latent_end>, and a placeholder token <latent_pad>. The reasoning segment is structurally defined as a sequence bounded by the start and end tokens, with the intermediate positions reserved for the latent embeddings. During Training, we replace the intermediate <latent_pad> tokens with the ground-truth latent sequence $\mathcal{Z}_{\text{GT}}$. This allows the model to learn the transition dynamics via standard teacher forcing. While during inference, the model is initialized with <latent_start> followed by a sequence of <latent_pad> tokens. The slow reason expert then autoregressively generates latent embeddings, sequentially filling the positions of the <latent_pad> placeholders until the pre-defined horizon is filled. Horizontal length can be adaptively adjusted and its effect is further analyzed through ablation studies.
Latent Supervision Strategy. Although inference is performed in an autoregressive manner, we train the slow reasoning expert using continuous latent regression rather than discrete token likelihoods (Yang et al., 2025; Wang et al., 2025b). Specifically, the slow expert is trained to predict a sequence of latent reasoning states $\hat{\mathcal{Z}}$ in a next-step prediction manner, conditioned on preceding observations and context. Unlike conventional CoT supervision based on discrete token prediction, our latent targets consist of continuous, high-dimensional embeddings that encode future physical world states. To align the predicted latent with the ground-truth representations, we employ cosine similarity as the supervision objective. The loss is defined as:
@@
By maximizing directional alignment in the latent space, this objective encourages the model to anticipate future physical dynamics in a structured and compact manner.
3.4 Dual-System Coordination
@@
Figure 3: The reasoning expert performs low-frequency latent CoT reasoning to capture spatio-temporal dependencies, while the fast acting expert generates actions conditioned on high-frequency observations and periodically updated latent knowledge.
@@
Asynchronous Frequency Coordination. To coordinate LaST CoT reasoning with high-frequency robotic control, we introduce an asynchronous frequency mechanism between the slow reasoning expert and the fast acting expert. As shown in Fig. 3, we decouple their operating frequencies using a set of update ratios $\kappa$ (e.g., $\kappa\in{2,4,8}$). The slow expert is activated only at sparse keyframes ($t\bmod\kappa=0$), where it performs autoregressive latent CoT reasoning. In contrast, the fast expert runs at the native control frequency and remains active at every timestep. Between consecutive keyframes ($t\bmod\kappa\neq 0$), the slow reasoning expert is held dormant, while the fast expert generates actions by conditioning on the most recent latent reasoning output. An interesting empirical finding is that increasing the temporal extent of the latent representations (e.g., by predicting multiple future keyframes) leads to improved performance in action prediction. For the inputs to the two experts, the slow reasoning expert receives the natural language instruction $l$ and the low-frequency observation $I_{\text{slow}}$, constructing the Latent CoT that encapsulates future physical dynamics. Conversely, the fast acting expert is optimized for rapid closed-loop feedback and receives only the high-frequency observation $I_{\text{fast}}$. Crucially, since our MoT architecture maintains a unified token sequence, the fast expert can efficiently attend to both the linguistic goal and the Latent CoT tokens via the shared attention mechanism. The description of the KV cache is provided in Appendix F.2.
3.5 Training Recipe
Large-Scale Robotic Pretraining. First, LaST0 performs pretraining on a diverse corpus of over 400K trajectories aggregated from Open-X-Embodiment (Open X-Embodiment Collaboration et al., 2023), DROID (Khazatsky et al., 2024), ROBOMIND (Wu et al., 2025), and other robotic datasets. Details are provided in Appendix A.
@@ 3.5 Training Recipe
Supervised Fine-Tuning (SFT). We employ a joint SFT strategy that optimizes both the slow reasoning expert and the fast action expert. Specifically, the slow reasoning expert is trained by minimizing the Latent CoT regression loss $\mathcal{L}_{\text{latent}}$, aligning its latent representations with domain-specific physical dynamics. In parallel, the fast acting expert is optimized using the standard Flow Matching loss $\mathcal{L}_{\text{flow}}$ for action denoising. Meanwhile, the action expert is trained with randomly mixed fast-slow operating ratios (e.g., 1:1, 1:2, 1:4), which exposes it to latent conditions updated at varying delays. As a result, LaST0 can flexibly accommodate different reasoning-update frequencies at deployment and adaptively choose the fast-slow inference rate. In ablation studies, we find that training with mixed fast-slow operating ratios does not degrade performance; instead, it improves the model’s robustness during inference.
@@
Table 1: Comparison of LaST0 and baselines on RLBench. All methods are trained in the multi-task setting (Shridhar et al., 2022), and we report mean success rates (S.R.). Inference speed is evaluated on an NVIDIA 4090 GPU.
@@
box laptop lid seat down to dustpan fridge on base out off hanger rack plants & Var $\downarrow$ speed $\uparrow$
OpenVLA 0.60 0.35 0.75 0.55 0.85 0.20 0.30 0.15 0.20 0.05 0.40 $\pm$0.02 6.3 Hz
SpatialVLA 0.80 0.70 0.85 0.20 0.80 0.15 0.25 0.40 0.15 0.30 0.46 $\pm$0.03 7.9 Hz
CogACT 0.90 0.80 0.95 0.50 0.85 0.50 0.55 0.45 0.30 0.25 0.61 $\pm 0.04$ 9.8 Hz
CoT-VLA 0.95 0.75 1.00 0.80 0.65 0.50 0.40 0.50 0.55 0.50 0.66 $\pm 0.03$ 1.1 Hz
$\pi_{0.5}$ 0.90 0.95 0.85 0.75 1.00 0.05 0.10 0.80 0.75 0.35 0.65 $\pm$0.04 13.8 Hz
HybridVLA 0.85 0.95 1.00 0.90 1.00 0.50 0.50 0.70 0.50 0.50 0.74 $\pm$0.04 6.1 Hz
LaST0 0.95 0.95 1.00 0.80 0.85 0.75 0.75 0.70 0.85 0.60 0.82 $\pm$0.03 15.4 Hz
@@
4 Experiment
Section 4.1 evaluates the manipulation performance and inference efficiency of LaST0 in simulation, while Section 4.2 conducts the ablation study of each component. Section 4.3 reports results on real-world tasks.
Figure 5: Ablation study on key design choices of LaST0. We analyze (a) the importance of different latent modalities, (b) the number of tokens allocated per latent modality, (c) the temporal coverage in latent reasoning, and (d) the collaboration frequency between reasoning and action experts. Results are reported as average success rates across 10 RLBench tasks.
4.1 Simulation Experiment
Data collection. We evaluate LaST0 on a diverse set of 10 tasks from the RLBench (James et al., 2020) benchmark, conducted in the CoppeliaSim simulation environment. All tasks are executed using a Franka Panda robotic arm with a single front-view observation. Demonstration data are collected by following pre-defined waypoints with motion planner (Sucan et al., 2012). Following the frame-sampling protocol adopted in (Shridhar et al., 2022), we construct a training dataset comprising 100 trajectories per task.
Training and evaluation protocol. We benchmark LaST0 against six representative state-of-the-art VLA models: OpenVLA (Kim et al., 2024), $\pi_{0.5}$ (Intelligence et al., 2025), CogACT (Li et al., 2024b), SpatialVLA (Qu et al., 2025), CoT-VLA (Zhao et al., 2025) and HybridVLA (Liu et al., 2025a). For all baselines, we initialize from the officially released pretrained checkpoints and follow the full fine-tuning configurations recommended by the paper. For CoT-VLA, we reimplement the method on top of Janus-Pro (Chen et al., 2025b), the same foundation model used in our approach, and reproduce its explicit CoT reasoning to ensure a fair comparison. For LaST0, each observation consists of a single RGB image resized to $384\times 384$, a point cloud uniformly subsampled to 1024 points, a language instruction obtained directly from the simulator, and the robot proprioceptive state synchronized with the predicted action. The model is trained for 300 epochs during the SFT stage, using the AdamW optimizer (Loshchilov and Hutter, 2017) across 8 NVIDIA A800 GPUs. Evaluation follows the protocol of prior work (Goyal et al., 2023). For each task, we perform 20 rollout trials using the final checkpoint, repeat the evaluation three times with different random seeds, and report the mean success rate together with its variance.
Quantitative and qualitative analysis. In Table 1, LaST0-3.3B achieves a mean success rate of 82% across 10 RLBench manipulation tasks. LaST0 is trained under a mixed fast-slow operating frequency and evaluated at a 1:4 ratio. In particular, LaST0 surpasses the strongest existing methods, HybridVLA-7B (74%), $\pi_{0.5}$-3B (65%), and CogACT-7B (61%), by margins of 8%, 17%, and 21%, respectively. Beyond the overall average, LaST0 attains the highest success rate on 7 out of 10 tasks, indicating consistent performance gains across diverse manipulation skills. These improvements primarily stem from the latent reasoning representations produced by the reasoning expert, which condition the action expert with compact latent states encoding future visual dynamics, 3D spatial structure, and robot proprioceptive information, enabling more stable and temporally coherent action generation. In terms of execution efficiency (without action chunking scheme), LaST0 operates at an inference speed of 15.4 Hz, which is significantly faster than explicit CoT methods (CoT-VLA: 1.1 Hz) and remains competitive with $\pi_{0.5}$ (13.8 Hz). As shown in Fig. 4, we compare the attention heatmaps of LaST0 against variants without CoT and with explicit CoT (CoT-VLA). While the no-CoT variant and the explicit CoT method fail to aggregate features over the manipulated objects and the robot, LaST0 exhibits a highly concentrated attention pattern, highlighting its superior spatio-temporal understanding.
4.2 Ablation Study
To validate the key design choices of LaST0, we conduct comprehensive ablation experiments on 10 RLBench tasks.
Importance of latent CoT modalities. As shown in Fig. 5(a), we assess each latent modality by ablating individual components while keeping all other parameters fixed at their optimal settings (i.e., latent tokens = 1, temporal coverage = 4). When using only the image, point cloud, or robot state latent, the model achieves success rates of 74%, 76%, and 75%, respectively, indicating that each modality-specific latent provides a strong basis for action generation. The combination of multiple modality latents continues to provide additional performance improvements, even when the manipulation accuracy is already high. These results validate the importance of modeling comprehensive physical dynamics in the latent space, and further demonstrate that enabling the model to autonomously reason about the relationship between the robot and its interactive environment is effective for robotic manipulation.
@@ 4.2 Ablation Study
Number of tokens per latent modality. As shown in Fig. 5 b), we study how the token budget allocated to each latent modality affects performance by varying the number of latent tokens while keeping other components fixed. When no latent token is used, performance drops substantially to 68%, indicating limited reasoning capacity. Introducing a single token for each modality leads to a sharp improvement (up to 82%), demonstrating that even a minimal latent representation is sufficient to establish an effective latent decision state. Further increasing the number of tokens yields no significant accuracy improvement, suggesting that high-level latent tokens can compactly encode physical information sufficient for effective reasoning.
@@
Temporal coverage in latent reasoning. As shown in Fig. 5 c), we investigate the effect of the temporal horizon used in latent reasoning by varying the number of future time steps encoded into the latent state. Performance improves consistently as the latent temporal coverage increases (from 68% to 82% when extending from 0 to 4 steps), indicating that incorporating longer temporal dependencies enables more informed latent decision states. Since adding further coverage beyond 4 steps does not significantly improve performance, we chose 4 steps as the final latent temporal coverage. The results suggest that extending the temporal horizon of latent prediction beyond that of action prediction is sufficient for improving action generation robustness. Due to our fast-slow system design, extending the temporal horizon of the latent space does not significantly affect action generation speed.
@@
Collaboration frequency between reasoning and action experts. As shown in Fig. 5 d), varying the collaboration frequency has a clear impact on task success. A series of ratios such as 1:1, 1:2, and 1:4 achieve comparable performance (75-79%), while overly infrequent collaboration (1:8) leads to a relative drop to 74%. Finally, the mixed strategy trained by combining data from all collaboration ratios achieves the best performance (82%), using a 1:4 ratio during testing. These results indicate that our frequency joint training scheme enables more robust coordination between the reasoning and action experts across tasks.
Table 2: Comparison across real-world manipulation tasks. We report success rates (S.R.) for standard single-arm and dual-arm tasks (Franka), mobile manipulation (AgileX), and dexterous manipulation (TienKung). Mean S.R. denotes the average success rate.
Models Wipe Press Place dish Place egg Scoop Open pot Mean Place egg on bread Arrange Sort Open Place
whiteboard stamp on rack on bread popcorn pick corn S.R. $\uparrow$ Step 1 Step 2 Step 3 dishes spoon drawer button
SpatialVLA 0.60 0.67 0.30 0.20 0.27 0.40 0.41 0.20 0.07 0.00
$\pi_{0.5}$ 0.60 0.73 0.60 0.47 0.53 0.60 0.59 0.47 0.20 0.07 0.47 0.20 0.67 0.53
CoT-VLA 0.53 0.60 0.66 0.33 0.33 0.53 0.50 0.33 0.13 0.07 0.33 0.13 0.53 0.40
LaST0 0.73 0.93 0.80 0.66 0.66 0.53 0.72 0.66 0.47 0.33 0.67 0.27 0.87 0.60
Embodiment Franka Emika Panda AgileX Mobile Tienkung Dex
4.3 Real-World Experiment